# Persistence of AR($1$) sequences with Rademacher innovations and linear mod $1$ transforms

### v2

@@ 3. Existence of a positive solution to equation (11)
4. The tail asymptotics
We first give a simple representation of the action of the killed transition operator $P$ on the space $U$ of functions defined in (18). To this end, let us introduce the following notation.
Denote by $\operatorname*{Var}[f]$ the total variation of a function $f$ on $[0,1/(1-a)]$. Then
@@ 4. The tail asymptotics
where we put $f(0-)=0$, is a bijective isometry between $(U,\|\cdot\|)$ and $(\mathbb{R}^{\varkappa_{a}^{\prime}+1},{\|\cdot\|}_{1})$.
@@
The main difficulty is when the orbit of zero under $T_{a}$ is infinite and aperiodic. In this case the operator $A$ acts on the space of infinite summable sequences $\ell_{1}$, and the adjoint operator $A^{*}$ on $\ell_{\infty}$ is defined by the standard duality (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) between $\ell_{1}$ and $\ell_{\infty}$. If $\varkappa_{a}^{\prime}<\infty$, we use (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) to denote the scalar product on $\mathbb{R}^{\varkappa_{a}^{\prime}+1}$.
@@
If the orbit is aperiodic (i.e. $\varkappa_{a}=\varkappa_{a}^{\prime}$), then the restriction of the killed transition operator $P$ to $U$ is equivalent to A (i.e. $P=M^{-1}AM$). The quantity $\lambda_{a}$, defined in (11), is the leading eigenvalue of $A$, and $v$ and $v^{*}$ are the eigenvectors of $A$ and $A^{*}$ corresponding to $\lambda_{a}$, respectively. Moreover, there exist constants $C_{1}>0$ and $\gamma\in(0,1)$ such that in either case we have
$\left\|\lambda_{a}^{-n}A^{n}u-\frac{(u,v^{*})}{(v,v^{*})}v\right\|_{1}\leq C_{1}\gamma^{n}\|u\|_{1}$ (30)
for every $u\in(\mathbb{R}^{\varkappa_{a}^{\prime}+1},{\|\cdot\|}_{1})$ and integer $n\geq 0$.
If the orbit is eventually periodic, then the restriction of $P$ to $U$ is equivalent to the linear operator $\hat{A}$ on $\mathbb{R}^{\varkappa_{a}^{\prime}+1}$ defined by
${(\hat{A}u)}_{k}={(Au)}_{k}+c_{\varkappa_{a}^{\prime}}u_{\varkappa_{a}^{\prime}}\mathbbm{1}\{k=k_{0}\},\qquad 0\leq k\leq\varkappa_{a}^{\prime},$
where $k_{0}$ is the smallest integer such that $T_{a}^{k_{0}}(0)=T_{a}^{\varkappa_{a}^{\prime}+1}(0)$. The operator $\hat{A}$ is irreducible. The leading eigenvalue of $\hat{A}$ is $\lambda_{a}$ and the corresponding right and left eigenvectors are $MV$ and $v^{*}$. Relation (30) holds with $A$ replaced by $\widehat{A}$.
@@
Recall that a bounded linear operator $Q$ on a Banach space is called quasi-compact if there exists a compact operator $Q_{c}$ such that $\rho(Q_{c})>\rho(Q-Q_{c})$, where $\rho$ stands for the spectral radius. Assertion (30) implies that $A$ is quasi-compact when the orbit of zero is aperiodic. This implies that the operator $P$ is quasi-compact on $U$ for all $a\in(\frac{1}{2},\frac{2}{3}]$.
We now use Proposition 12 to establish tail asymptotics (10) with the constant given by
@@
Notably, for $a=2/3$ and $p=1/2$, the function $cV$ is a probability density by (9), (66), and Eq. (4.2) by Parry [18].
We first assume that the orbit of zero is aperiodic, i.e. $\varkappa_{a}=\varkappa_{a}^{\prime}$. From (30) we have
$A^{n}u\sim\frac{(u,v^{*})}{(v,v^{*})}\lambda_{a}^{n}v\quad\text{as }n\to\infty$ (32)
for every vector $u\in(\mathbb{R}^{\varkappa_{a}^{\prime}},{\|\cdot\|}_{1})$. We compute the constant factor using the equalities $\varkappa_{a}=\varkappa_{a}^{\prime}$ and (11):
@@
Now assume that the orbit of zero is eventually periodic. The Perron–Frobenius theorem applies to $\hat{A}$, therefore relations (30) and (32) remain valid if we substitute $A$ by $\hat{A}$ and $v$ by $MV$. To show that the constant factor is again $1/c$, we put $n_{0}=\varkappa^{\prime}-k_{0}+1$ and let $\bar{v}_{k}^{*}$ be given by the r.h.s. of (29) for all $k\geq 0$. The sequence $\{\delta_{k}\}$ is eventually periodic, hence so is $\{L_{k}\}$, and hence so is $\{\bar{v}_{k}^{*}\}$. Therefore, using equality (49),
@@
To state the above results in terms of the operator $P$, use that $P^{n}f=M^{-1}A^{n}Mf$. We have $M^{-1}v=V$ if the orbit of zero is aperiodic, and $M^{-1}(MV)=V$ otherwise. Therefore, for every $a\in(\frac{1}{2},\frac{2}{3}]$, we have
$P^{n}f(x)$ $\sim c(Mf,v^{*})\lambda_{a}^{n}V(x),\quad f\in U,$ (33)
where the rate of convergence is estimated by
$\|\lambda_{a}^{-n}P^{n}f-c(Mf,v^{*})V(x)\|\leq C_{1}\gamma^{n}\|f\|,\quad f\in U.$ (34)
Clearly, $\lambda_{a}$ is the leading eigenvalue of $P$ and the corresponding eigenfunction is $V$.
Finally, we determine the tail asymptotics for $\tau$. Since $\mathbb{P}_{x}(\tau>n)=P^{n}\mathbbm{1}$ and the constant function $\mathbbm{1}(x)\equiv 1$ belongs to $U$ and $M\mathbbm{1}=(1,0,0,\ldots)$, we get $(M\mathbbm{1},v^{*})=1$ and asymptotic relation (10) follows from (33).
@@
$\sum_{k=0}^{\varkappa-1}u_{k}{\mathbbm{1}}\{x\geq(T^{k}(0)-1)/a\}=\sum_{k=0}^{\varkappa}u_{k}\delta_{k}+\sum_{k=0}^{\varkappa-1}u_{k}(1-\delta_{k}){\mathbbm{1}}\{x\geq T^{k+1}(0)\}.$ (35)
@@
$\text{if }\ T^{k}(0)\geq 1,\ \text{ then }\ (T^{k}(0)+1)/a>1/(1-a)\ \text{ and}\ {\mathbbm{1}}\{x\geq(T^{k}(0)+1)/a\}=0$
because in the case $a=2/3$, where $2/a=1/(1-a)$, we have $T^{k}(0)\neq 1$; see Section 2.2. If $\varkappa$ is finite, then $T^{\varkappa}(0)\in(\frac{2a-1}{1-a},1)$ and therefore, $(T^{\varkappa}(0)+1)/a>1/(1-a)$. Hence
@@
$\begin{array}[]{rlc}&\lambda_{a}u_{k}^{*}=p\delta_{k}u_{0}^{*}+c_{k}u^{*}_{k+1},&0\leq k<\varkappa,\\ &\lambda_{a}u^{*}_{\varkappa}=p\delta_{\varkappa}u_{0}^{*}&\text{if }\varkappa<\infty.\\ \end{array}$ (41)
@@
Notice that in the case of infinite $\varkappa$, the vectors $u^{*}$ satisfy $u^{*}\in\ell_{\infty}$ by $\lambda_{a}>p$ if $p\geq 1/2$ and by (23) if $p<1/2$. In the case of finite $\varkappa$, the equation in the second line of (41) is satisfied since $\delta_{\varkappa}=1$. It follows that the vector $v^{*}$, which is proportional to such $u^{*}$, is the eigenvector of $A$ corresponding to $\lambda_{a}$.
@@
This implies that $R(|\lambda|)>1$, hence $|\lambda|<\lambda_{a}$. Thus, $\lambda_{a}$ is the leading eigenvalue of $A$ when $\varkappa$ is finite, which concludes the proof in this case. From now on we assume that the orbit of zero is infinite (and still aperiodic).
Next we notice that all eigenvalues of $A$ other than $\lambda_{a}$ are contained in the centred ball of radius $\lambda_{a}-\varepsilon$ with some $\varepsilon>0$. Otherwise there is a convergent sequence of eigenvalues $\{\mu_{n}\}$ such that $|\mu_{n}|\to|\lambda_{a}|$. If $p\geq 1/2$, then $R$ is analytic on the set $D=\{\lambda\in\mathbb{C}:|\lambda|>p\}$, which contains $\mu_{n}$ for all $n$ large enough by $\lambda_{a}>p$. If $p<1/2$, then it follows from (23) that $R$ is analytic on the set $D=\{\lambda\in\mathbb{C}:|\lambda|>p(q/p)^{C}\}$, which contains $\mu_{n}$ for all $n$ large enough since $\lambda_{a}>p(q/p)^{C}$ by (22). In either case it must be that $R\equiv 1$ on $D$ by $R(\mu_{n})\equiv 1$, which is a contradiction.
We now show that the spectrum of $A$ without the point $\lambda_{a}$ is contained in the centred ball of radius $\lambda_{a}-\varepsilon$. Equivalently, the resolvent operator $(A-\lambda I)^{-1}$ is bounded for every $\lambda\neq\lambda_{a}$ with $|\lambda|>\lambda_{a}-\varepsilon$.
@@
is the unique solution to the equations. Therefore,
@@
and thus $u_{0}$ is defined uniquely.
@@
$\left|\lambda(R(\lambda)-1)\right||u_{0}|$ $\leq|w_{0}|+p\sum_{k=1}^{\infty}\delta_{k}\sum_{j=1}^{k}|w_{j}|\frac{p^{k-j}}{|\lambda|^{k-j+1}}$
$\leq|w_{0}|+\frac{p}{|\lambda|-p}\sum_{j=1}^{\infty}|w_{j}|.$ (47)
Consequently, for every $\lambda\neq\lambda_{a}$ with $|\lambda|>\lambda_{a}-\varepsilon$ there exists a constant $C(\lambda)$ such that
@@
We know from (22) that $\lambda_{a}>pr^{C}$. Hence there exists an $\varepsilon>0$ such that (48) is also valid in the case $p<1/2$. Consequently, for every $p\in(0,1)$, the spectrum of $A$ without the point $\lambda_{a}$ is contained in the centred ball of radius $\lambda_{a}-\varepsilon$.
The property of the spectrum shown just above implies that the operator $A$ is quasi-compact. If one shows additionally that $A$ possesses certain additional positivity properties, then one can apply a Krein–Rutman type result, for example Theorems 6 and 7 in Sasser [21]. In our particular case we will arrive at the same conclusion by direct calculations, which are quite standard for the area of quasi-compact operators.
Define the linear operators $Qu=\frac{(u,v^{*})}{(v,v^{*})}v$ and $Bu=A(u-Qu)$ on $\ell_{1}$. Clearly, $Q^{2}=Q$. Since $v$ and $v^{*}$ are the eigenvectors corresponding to $\lambda_{a}$, we have $AQu=\lambda_{a}Qu$ and
@@
Thus, $A$ and $Q$ are commuting and $QBu\equiv 0$. These properties of the operators $A$, $B$, and $Q$ imply that
$A^{n}u=\lambda_{a}^{n}Qu+B^{n}u,\quad u\in\ell_{1},\ n\geq 1.$
Thus, (30) will follow from this representation if we show that the spectral radius of $B$ is strictly smaller than $\lambda_{a}$. To this end we consider the closed linear subspace
@@
Since $Bu\in Y$ for any $u\in\ell_{1}$, it suffices to consider the restriction of $B$ to this subspace. We first show that the spectrum of $B$ is a subset of the spectrum of $A$. Let $\lambda$ be any regular value of $A$, i.e. $(A-\lambda I)u=w$ is solvable for every $w\in\ell_{1}$. Assume now that $w\in Y$ and let $u_{w}$ denote the solution to $(A-\lambda I)u=w$. Then
00 0 $=(w,v^{*})=((A-\lambda I)u_{w},v^{*})=(Au_{w},v^{*})-\lambda(u_{w},v^{*})$
@@
$(B-\lambda I_{Y})u_{w}=(A-\lambda I)u_{w}=w$
and $\lambda$ is regular for $B$. Since the spectrum of $A$ without point $\lambda_{a}$ is contained in a centred ball of radius $\lambda_{a}-\varepsilon$, it remains to show that $\lambda_{a}$ is regular for $B$. To solve $(B-\lambda_{a}I_{Y})u=w$, it suffices to take $\lambda=\lambda_{a}$ in (44) and determine $u_{0}$ from the equation $(u,v^{*})=0$. Alternatively, we can argue as follows.
Let $u^{(\lambda)}$ denote the solution to $(B-\lambda I_{Y})u=w$ for $|\lambda-\lambda_{a}|<\delta$ with some sufficiently small $\delta>0$. Then, by (45),
@@
Using now (44) with $\lambda=\lambda_{a}$ and $u_{0}^{(\lambda_{a})}$ defined just above, we find the unique solution $u^{(\lambda_{a})}$ to $(B-\lambda_{a}I_{Y})u=w$. It follows from Proposition 8 that $u^{(\lambda_{a})}\in\ell_{1}$. This completes the proof in the aperiodic case.
@@
$\begin{array}[]{cl}&(\widehat{A}s)_{0}=p\sum_{k=0}^{\varkappa^{\prime}}\delta_{k}s_{k}+c_{\varkappa^{\prime}}s_{\varkappa^{\prime}}\mathbbm{1}\{k_{0}=0\},\\ &(\widehat{A}s)_{k}=c_{k-1}s_{k-1}+c_{\varkappa^{\prime}}s_{\varkappa^{\prime}}\mathbbm{1}\{k_{0}=k\},\quad 1\leq k\leq\varkappa^{\prime}.\\ \end{array}$ (50)
We first note that the operator $\hat{A}$ is irreducible, meaning that for every $0\leq i,j\leq\varkappa^{\prime}$ there exists an $n=n(i,j)$ such that the matrix representing $\widehat{A}^{n}$ has a strictly positive entry at the position $(i,j)$. Combining this with the fact that the $(0,0)$-entry in the matrix corresponding to $\widehat{A}$ equals $p>0$, we conclude that there exists an $n_{0}\geq 1$ such that the matrix representing $\widehat{A}^{n_{0}}$ is strictly positive. In other words, $\widehat{A}$ is aperiodic. Therefore, the classical Perron–Frobenius theorem is applicable, and $\widehat{A}$ has a positive eigenvalue which strictly exceeds the absolute values of all other eigenvalues.
@@
The positive solution to $\lambda^{\varkappa^{\prime}-k_{0}+1}=c_{k_{0}}c_{k_{0}+1}\cdots c_{\varkappa^{\prime}}$ is not an eigenvalue of $\hat{A}$. Indeed, in this case $s_{0}=0$, and plugging (51) and (52) into the first line of (50) gives an impossible identity
@@
We now easily see that (56) is equivalent to (11) in the purely periodic case. Therefore, $\lambda_{a}$ is the leading eigenvalue of $\hat{A}$ for all possible $k_{0}$.
@@
Lastly, we find the left eigenvectors $u^{*}\hat{A}=\lambda_{a}u^{*}$. Using the definition of $\hat{A}$, we get the coordinate-wise equations
$\begin{array}[]{rlc}&\lambda_{a}u_{k}^{*}=p\delta_{k}u_{0}^{*}+c_{k}u^{*}_{k+1},&0\leq k<\varkappa^{\prime},\\ &\lambda_{a}u^{*}_{\varkappa^{\prime}}=p\delta_{\varkappa^{\prime}}u_{0}^{*}+c_{\varkappa^{\prime}}u^{*}_{k_{0}}.\\ \end{array}$
We solve them exactly as we did in the aperiodic case. This gives us the same answer $u_{k}^{*}=v_{k}^{*}$ for $1\leq k\leq\varkappa^{\prime}$ because the first $\varkappa^{\prime}$ equations are the same and we did not use the equation in the second line. It is not hard to check that this equation is indeed satisfied for such $u^{*}$. Then the vector $v^{*}$, which is proportional to $u^{*}$, is the left eigenvector of $\hat{A}$ corresponding to $\lambda_{a}$.

### v3

@@ 3. Existence of a positive solution to equation (15)
4. Spectral properties of the killed transition operator on $U$
In this section we establish the tail asympotics of $\tau$, given in (15). To do this, we first provide a simple representation of the action of the killed transition operator $P$ on the space $U$ of functions defined in (28). To this end, let us introduce the following notation.
Denote by $\operatorname*{Var}[f]$ the total variation of a function $f$ on [0,1/(1a)][0,1/(1-a)]. Then
@@ 4. Spectral properties of the killed transition operator on $U$
where we put f(0)=0f(0-)=0, is a bijective isometry between $(U,\|\cdot\|)$ and $(\mathbb{R}^{\varkappa_{a}^{\prime}+1},{\|\cdot\|}_{1})$.
@@
The main difficulty is when the orbit of zero under $T_{a}$ is infinite and aperiodic. In this case the operator $A$ acts on the space of infinite summable sequences $\ell_{1}$, and the adjoint operator $A^{*}$ on $\ell_{\infty}$ is defined by the standard duality $(\cdot,\cdot)$ between $\ell_{1}$ and $\ell_{\infty}$. If $\varkappa_{a}^{\prime}<\infty$, we use $(\cdot,\cdot)$ to denote the scalar product on $\mathbb{R}^{\varkappa_{a}^{\prime}+1}$.
@@
If the orbit is aperiodic (i.e. $\varkappa_{a}=\varkappa_{a}^{\prime}$), then the restriction of the killed transition operator $P$ to $U$ is equivalent to A (i.e. $P=M^{-1}AM$). The quantity $\lambda_{a}$, defined in (15), is the leading eigenvalue of $A$ (i.e. every other eigenvalue $\lambda$ satisfies $|\lambda|<\lambda_{a}$), and $v$ and $v^{*}$ are the eigenvectors of $A$ and $A^{*}$ corresponding to $\lambda_{a}$, respectively. Moreover, there exist constants $C_{1}>0$ and $\gamma\in(0,1)$ such that for every $u\in(\mathbb{R}^{\varkappa_{a}^{\prime}+1},{\|\cdot\|}_{1})$ and integer $n\geq 0$,
$\left\|\lambda_{a}^{-n}A^{n}u-\frac{(u,v^{*})}{(v,v^{*})}v\right\|_{1}\leq C_{1}\gamma^{n}{\|u\|}_{1}.$ (39)
If the orbit is eventually periodic, then the restriction of $P$ to $U$ is equivalent to the linear operator $\widehat{A}$ on $\mathbb{R}^{\varkappa_{a}^{\prime}+1}$ defined by
${(\widehat{A}u)}_{k}={(Au)}_{k}+c_{\varkappa_{a}^{\prime}}u_{\varkappa_{a}^{\prime}}\mathbbm{1}\{k=k_{0}\},\qquad 0\leq k\leq\varkappa_{a}^{\prime},$
where $k_{0}$ is the smallest integer such that $T_{a}^{k_{0}}(0)=T_{a}^{\varkappa_{a}^{\prime}+1}(0)$. Then $\lambda_{a}$ is leading eigenvalue of $\widehat{A}$, and $Mv$ and $v^{*}$ are the eigenvectors of $\widehat{A}$ and $^{\widehat{A}}^{*}$ corresponding to $\lambda_{a}$, respectively. Relation (39) remains true when $A$ is replaced by $\widehat{A}$ and $v$ is replaced by $MV$.
@@
Recall that a bounded linear operator $Q$ on a Banach space is called quasi-compact if there exists a compact operator $Q_{c}$ such that $\rho(Q)>\rho(Q-Q_{c})$, where $\rho$ stands for the spectral radius. Assertion (39) implies that $A$ and $\widehat{A}$ are quasi-compact. This implies that the operator $P$ is quasi-compact on $U$ for all $a\in(\frac{1}{2},\frac{2}{3}]$.
We postpone the proof of Proposition 14 and use this result to establish the tail asymptotics (14) with the constant
@@
We first assume that the orbit of zero is aperiodic, i.e. $\varkappa_{a}=\varkappa_{a}^{\prime}$. We compute the constant factor $(v,v^{*})$ in (39) using the equality $\varkappa_{a}=\varkappa_{a}^{\prime}$ and the definition (15) of $\lambda_{a}$:
@@
Now assume that the orbit of zero is eventually periodic. Relation (39) is valid with $A$ replaced by $\widehat{A}$ and $v$ replaced by $MV$. Let us compute the constant factor $(Mv,v^{*})$ there. Put $n_{0}=\varkappa^{\prime}-k_{0}+1$ and let $\bar{v}_{k}^{*}$ be defined as in (38) for all $k\geq 0$. The sequence $\{\delta_{k}\}$ is eventually periodic, hence so is $\{L_{k}\}$, and hence so is $\{\bar{v}_{k}^{*}\}$. Therefore, using equality (59),
@@
To state (39) in terms of the operator $P$, use that $P^{n}f=M^{-1}A^{n}Mf$. We have $M^{-1}v=V$ if the orbit of zero is aperiodic, and $M^{-1}(MV)=V$ if it is eventually periodic. Therefore, (39) directly implies that for every $a\in(\frac{1}{2},\frac{2}{3}]$, we have
$\|\lambda_{a}^{-n}P^{n}f-c(Mf,v^{*})V\|\leq C_{1}\gamma^{n}\|f\|,\quad f\in U.$ (41)
Clearly, $\lambda_{a}$ is the leading eigenvalue of $P$ and the corresponding eigenfunction is $V$. Since
$|f(x)|\leq|f(0)|+|f(x)-f(0)|\leq\|f\|$ (42)
for every $x\in[0,\tfrac{1}{1-a}]$, it follows from (41) that
$\sup_{0\leq x\leq\frac{1}{1-a}}\Big|\lambda_{a}^{-n}P^{n}f(x)-c(Mf,v^{*})V(x)\Big|\leq C_{1}\gamma^{n}\|f\|,\quad f\in U.$ (43)
Finally, we determine the tail asymptotics for $\tau$. We have $\mathbb{P}_{x}(\tau>n)=P^{n}\mathbbm{1}(x)$ and the constant function $\mathbbm{1}(x)\equiv 1$ belongs to $U$. Since $M\mathbbm{1}=(1,0,0,\ldots)$, we get $(M\mathbbm{1},v^{*})=1$ by (38), and the asymptotic relation (14) follows from (43).
@@
$\sum_{k=0}^{\varkappa}u_{k}{\mathbbm{1}}\{x\geq(T^{k}(0)-1)/a\}=\sum_{k=0}^{\varkappa}u_{k}\delta_{k}+\sum_{k=0}^{\varkappa-1}u_{k}(1-\delta_{k}){\mathbbm{1}}\{x\geq T^{k+1}(0)\}.$ (44)
@@
$\text{if }\ T^{k}(0)\geq 1,\ \text{ then }\ (T^{k}(0)+1)/a>1/(1-a)\ \text{ and }\ {\mathbbm{1}}\{x\geq(T^{k}(0)+1)/a\}=0,$
because in the case $a=2/3$, where 2/a=1/(1a)2/a=1/(1-a), we have $T^{k}(0)\neq 1$; see Section 2.3. If $\varkappa$ is finite, then $T^{\varkappa}(0)\in(\frac{2a-1}{1-a},1)$ and therefore, $(T^{\varkappa}(0)+1)/a>1/(1-a)$. Hence
@@
[…]
@@
Notice that in the case of infinite $\varkappa$, the vectors $u^{*}$ satisfy $u^{*}\in\ell_{\infty}$ by $\lambda_{a}>p$ if $p\geq 1/2$ and by (32) if $p<1/2$. In the case of finite $\varkappa$, the equation in the second line of (50) is satisfied since $\delta_{\varkappa}=1$. It follows that the vector $v^{*}$, which is proportional to such $u^{*}$, is the eigenvector of $A^{*}$ corresponding to $\lambda_{a}$.
@@
This implies that $R(|\lambda|)>1$, hence $|\lambda|<\lambda_{a}$. Thus, $\lambda_{a}$ is the leading eigenvalue of $A$. Denote $\varrho^{\prime}=0$ if $A$ has no other eigenvalues, otherwise put
$\varrho^{\prime}=\sup\{|\lambda|:\lambda\text{ is an eigenvalue of }A,\lambda\neq\lambda_{a}\}.$
Assume that $\varkappa$ is finite. It is easy to see that the matrix $\mathcal{A}$ that represents $A$ in the standard basis, has non-negative entries and is irreducible, that is for some $m\in\mathbb{N}$, all entries of $\mathcal{A}+{\mathcal{A}}^{2}+\ldots+{\mathcal{A}}^{m}$ are strictly positive. Then $\lambda_{a}$ is a simple root of the characteristic polynomial of $\mathcal{A}$ and ${\mathcal{A}}^{n}/\lambda_{a}^{n}\to vv^{*}/v^{*}v$ (with $v^{*}$ regarded as a row vector) in the operator norm as $n\to\infty$ by the classical Perron–Frobenius theorem; see Meyer [19, p. 673 and Eq. (8.3.10)]. For the rate of this convergence, the spectral resolution theorem for $A^{n}$ implies the bound (39) for any $\gamma\in(\varrho^{\prime}/\lambda_{a},1)$; see Eq. (7.9.9) and the last formula on p. 629 in [19].
From now on we assume that the orbit of zero is infinite (and still aperiodic). Notice that $\varrho^{\prime}<\lambda_{a}$, otherwise there is a convergent sequence of eigenvalues $\{\mu_{n}\}$ such that $|\mu_{n}|\to|\lambda_{a}|$. If $p\geq 1/2$, then $R$ is analytic on the set $D=\{\lambda\in\mathbb{C}:|\lambda|>p\}$, which contains $\mu_{n}$ for all $n$ large enough by $\lambda_{a}>p$. If $p<1/2$, then it follows from (32) that $R$ is analytic on the set $D=\{\lambda\in\mathbb{C}:|\lambda|>p(q/p)^{C}\}$, which contains $\mu_{n}$ for all $n$ large enough since $\lambda_{a}>p(q/p)^{C}$ by (31). In either case it must be that $R\equiv 1$ on $D$ by $R(\mu_{n})\equiv 1$, which is a contradiction.
We now show that the spectrum of $A$ without the point $\lambda_{a}$ is contained in the closed centred ball of radius $\varrho=\max(\varrho^{\prime},p,p(q/p)^{C})$, where $\varrho<\lambda_{a}$. Equivalently, the resolvent operator $(A-\lambda I)^{-1}$ is bounded for every $\lambda\neq\lambda_{a}$ with $|\lambda|>\varrho$.
@@
is the unique solution to the equations in the second line of (52). Therefore,
@@
and thus $u_{0}$ is defined uniquely since $\lambda$ is not an eigenvalue of $A$, and therefore the factor of $u_{0}$ on the left-hand side is non-zero by (48).
@@
$\left|\lambda(R(\lambda)-1)\right||u_{0}|$ $\leq|w_{0}|+p\sum_{k=1}^{\infty}\delta_{k}\sum_{j=1}^{k}|w_{j}|\frac{p^{k-j}}{|\lambda|^{k-j+1}}\leq|w_{0}|+\frac{p}{|\lambda|-p}\sum_{j=1}^{\infty}|w_{j}|.$ (56)
Consequently, for every $\lambda\neq\lambda_{a}$ with $|\lambda|>\varrho$ there exists a constant $C(\lambda)$ such that
@@
We know from (31) that $\lambda_{a}>pr^{C}$, hence (57) is also valid in the case $p<1/2$. Consequently, for every $p\in(0,1)$, the spectrum of $A$ without the point $\lambda_{a}$ is contained in the closed centred ball of radius $\varrho<\lambda_{a}$.
The property of the spectrum shown just above implies that the operator $A$ is quasi-compact. If one shows additionally that $A$ possesses certain additional positivity properties, then one can apply a Krein–Rutman type result, for example Theorems 6 and 7 in Sasser [23]. In our particular case we will arrive at the same conclusion by the direct calculations presented below, which are quite standard for the area of quasi-compact operators.
Define the linear operators $Qu=\frac{(u,v^{*})}{(v,v^{*})}v$ and Bu=A(uQu)Bu=A(u-Qu) on $\ell_{1}$, where $(v,v^{*})\geq v_{0}v_{0}^{*}>0$ by (38). Clearly, $Q^{2}=Q$. Since $v$ and $v^{*}$ are the eigenvectors corresponding to $\lambda_{a}$, we have $AQu=\lambda_{a}Qu$ and
@@
Thus, $A$ and $Q$ are commuting and $QB=BQ=0$. These properties of the operators $A$, $B$, and $Q$ imply that
$A^{n}u=\lambda_{a}^{n}Qu+B^{n}u,\quad u\in\ell_{1},\ n\geq 1.$ (58)
We claim that the spectral radius $\rho(B)$ of $B$ satisfies $\rho(B)\leq\varrho$. Then the estimate (39) for any $\gamma\in(\varrho/\lambda_{a},1)$ follows follow from the representation (58) combined with Gelfand’s formula $\rho(B)=\lim_{n\to\infty}\|B^{n}\|^{1/n}$.
To prove the claim, we consider the closed linear subspace
@@
Since $Bu\in Y$ for any $u\in\ell_{1}$, it suffices to consider the restriction of $B$ to this subspace, which we denote by $B_{Y}$. In fact, we have $\rho(B)=\rho(B_{Y})$, where the inequality $\rho(B)\geq\rho(B_{Y})$ is immediate from Gelfand’s formula, and the reverse one follows from the said formula by
$\rho(B)=\lim_{n\to\infty}\sup_{u\in\ell_{1}:\|u\|\leq 1}\|B^{n}u\|^{1/n}\leq\lim_{n\to\infty}\sup_{v\in Y:\|v\|\leq\|B\|}\|B^{n-1}v\|^{1/n}=\rho(B_{Y}).$
We first show that the spectrum of $B_{Y}$ is a subset of the spectrum of $A$. Let $\lambda$ be any regular value of $A$, i.e. $(A-\lambda I)u=w$ is solvable for every $w\in\ell_{1}$. Assume now that $w\in Y$ and let $u_{w}$ denote the solution to $(A-\lambda I)u=w$. Then
$0$ $=(w,v^{*})=((A-\lambda I)u_{w},v^{*})=(Au_{w},v^{*})-\lambda(u_{w},v^{*})$
@@
$(B_{Y}-\lambda I_{Y})u_{w}=(A-\lambda I)u_{w}=w$
and $\lambda$ is regular for $B_{Y}$. Since the spectrum of $A$ without point $\lambda_{a}$ is contained in the closed centred ball of radius $\varrho$, it remains to show that $\lambda_{a}$ is regular for $B_{Y}$. We need to show that the equation $(B_{Y}-\lambda_{a}I_{Y})u=w$ for $w\in Y$ has a solution $u\in Y$. Since $B_{Y}=A$ on $Y$, this is equivalent to showing that $(A-\lambda_{a}I)u=w$ has a solution $u\in Y$. Therefore, we can take $\lambda=\lambda_{a}$ in (53) and notice that (54) is satisfied when $\lambda=\lambda_{a}$ for any $u_{0}$ because in this case the right-hand side of (54) is $(w,v^{*})$. It remains to determine $u_{0}$ from the equation $(u,v^{*})=0$.
Alternatively, we can solve $(B_{Y}-\lambda_{a}I_{Y})u=w$ as follows. Let $u^{(\lambda)}$ denote the solution to $(B_{Y}-\lambda I_{Y})u=w$ for $\lambda$ such that $0<|\lambda-\lambda_{a}|<\lambda_{a}-\varrho$. Then, by (54),
@@
Using now (53) with $\lambda=\lambda_{a}$ and $u_{0}=u_{0}^{(\lambda_{a})}$ defined just above, we find the unique solution $u^{(\lambda_{a})}$ to $(B_{Y}-\lambda_{a}I_{Y})u=w$. It follows from Proposition 10 that $u^{(\lambda_{a})}\in\ell_{1}$. Then $u^{(\lambda_{a})}\in Y$ since $Y$ is closed in $\ell_{1}$. This completes the proof in the aperiodic case.
@@
[…]
We first note that the operator $\widehat{A}$ is non-negative and irreducible. Moreover, the matrix that represents $\widehat{A}$ in the standard basis is primitive, i.e. it does not have a cyclic structure, see [19, p. 680]. Therefore, by the Perron–Frobenius theorem, $\widehat{A}$ has a positive simple eigenvalue that strictly exceeds the maximum $\varrho^{\prime}$ of absolute values of all other eigenvalues.
@@
The positive solution to $\lambda^{\varkappa^{\prime}-k_{0}+1}=c_{k_{0}}c_{k_{0}+1}\cdots c_{\varkappa^{\prime}}$ is not an eigenvalue of $\widehat{A}$. Indeed, in this case $s_{0}=0$, and plugging (61) and (62) into the first line of (60) gives an impossible identity
@@
We now easily see that (66) is equivalent to (15) in the purely periodic case. Therefore, $\lambda_{a}$ is the leading eigenvalue of $\widehat{A}$ for all possible $k_{0}$.
@@
Lastly, we find the eigenvectors ${\widehat{A}}^{*}u^{*}=\lambda_{a}u^{*}$. Using the definition of $\widehat{A}$, we get the coordinate-wise equations
[…]
We solve them exactly as we did in the aperiodic case for (50). This gives us the same answer $u_{k}^{*}=u_{0}^{*}v_{k}^{*}$ for $1\leq k\leq\varkappa^{\prime}$ because the first $\varkappa^{\prime}$ equations are the same and we did not use the equation in the second line. It is not hard to check that this last equation is indeed satisfied for such $u^{*}$. Then $v^{*}$, which is proportional to $u^{*}$, is the eigenvector of $^{\widehat{A}}^{*}$ corresponding to $\lambda_{a}$. Therefore, (39) follows for any $\gamma\in(\varrho^{\prime}/\lambda_{a},1)$ as in the aperiodic case with finite $\varkappa$.